OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log6(5x - 5) = log6(3x + 7) A. {1} B. {6} C. Ø D. {2}

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@rafabc02 @Romero Can someone plz help me!

OpenStudy (anonymous):

log6(5x - 5) = log6(3x + 7) As we have the same log both side we can set the "inside" equal: 5x-5=3x+7 How do we continue now?

OpenStudy (anonymous):

oh so u just solve for x now?

OpenStudy (anonymous):

so its B? @rafabc02

OpenStudy (anonymous):

Yes! Well done! @sarah43

OpenStudy (anonymous):

Any doubt? PD: Thanks for the Medal! :D

OpenStudy (anonymous):

nope. thnx alot!

OpenStudy (anonymous):

You deserve it!